# A Triple Scenario for the Formation of Wide Black Hole Binaries Such As Gaia BH1

### v1

@@ Abstract
Recently, several non-interacting black hole-stellar binaries have been identified in Gaia data. For example, Gaia BH1, where a Sun-like star is in moderate eccentricity (e=0.45), 185-day orbit around a black hole. This orbit is difficult to explain through binary evolution. The present-day separation suggests the progenitor system would have undergone an episode of common envelope evolution, but a common envelope should shrink the period below the observed one. Since the majority of massive stars form in higher multiplicity systems, a triple evolution scenario is more likely for the progenitors of BH binaries. Here we show that such systems can indeed be more easily explained via evolution in hierarchical triple systems. von Zeipel–Lidov–Kozai oscillations or instabilities can delay the onset of the common envelope phase in the inner binary of the triple, so that the black hole progenitor and low-mass star are more widely separated when it begins, leading to the formation of wider binaries. There are also systems with similar periods but larger eccentricities, where the BH progenitor is a merger product of the inner binary in the triple. Such mergers lead to a more top-heavy black hole mass function.
@@ 1 Introduction
Approximately one out of every thousand stars will end their lives as a black hole (BH). This means a galaxy like the Milky Way should contain of order black holes. However, only of order a hundred black hole systems are known in our Galaxy, predominantly from studies of X-ray binaries (Corral-Santana et al., 2016; Fortin et al., 2023), that can only identify close, interacting systems.
Microlensing (Lam et al., 2022; Mróz et al., 2022; Sahu et al., 2022), spectroscopy (Shenar et al., 2022), and astrometry (Andrews et al., 2022; Shahaf et al., 2023) can be used to identify isolated black holes or non-interacting binaries. Such systems probe high mass binary evolution over a wider range of parameter space than X-ray binaries, providing important constraints on the black hole natal kick distribution and common envelope physics.
Gaia DR3 contains a few binary orbital solutions (Gaia Collaboration et al., 2023a, b), prompting searches for astrometric BH binaries (Andrews et al., 2022; Shahaf et al., 2023; El-Badry et al., 2023a, b). To date, two unambiguous BH-stellar binaries have been identified: Gaia BH1 and Gaia BH2. Gaia BH1 contains a G-type main sequence star in a 185.6-day orbit around a black hole. The star’s orbit is moderately eccentric with (El-Badry et al., 2023a). Gaia BH2 contains a red giant in a 1276.7 day orbit around a black hole. This star’s orbit is also moderately eccentric with (El-Badry et al., 2023b). The uncertainties on the orbital periods and eccentricities are much less than 1%.
These systems pose a challenge to isolated binary evolution models. On the one hand, the present-day separations suggest the progenitor system would have undergone an episode of common envelope evolution. On the other hand, a common envelope would shrink the binary period below the observed one. The tension can be resolved by invoking an unusually large common envelope efficiency ( for Gaia BH1; El-Badry et al., 2023a), or a very massive progenitor for the black hole that avoids a red giant phase (and hence the common envelope) entirely (El-Badry et al., 2023b).
In any case, an isolated binary is not the most likely initial configuration for the progenitor system. Most O & B stars are in triples or higher order multiples (Moe & Di Stefano, 2017). Stellar evolution in triples could give rise to a variety of novel evolutionary channels and produce different types of binaries and merger products (e.g. Perets & Fabrycky, 2009; Perets & Kratter, 2012; Antognini et al., 2014; Michaely & Perets, 2014; Naoz, 2016; Antonini et al., 2017; Rose et al., 2019; Stegmann et al., 2022). Motivated by this observation we check whether triple evolution can reproduce the Gaia black hole binaries. We find that triples can reproduce the observed period and eccentricity of Gaia BH1, even for normal common envelope efficiencies ().
@@ 2.1 Initial conditions
The progenitor of the black hole in Gaia BH1 would be , while its companion is 0.93 . Thus, the initial mass ratio is . The distribution of binary properties for such low mass ratios is unconstrained by Moe & Di Stefano (2017). Here, we extrapolate the broken power-law mass ratio distribution from their work down to the brown dwarf boundary. This extrapolation affects the distribution of companion periods, as shown in Figure 1.
@@ 3 Results
The top, left panels of Figure 3 and Figure 4 show the masses and orbital eccentricities for surviving black hole-main sequence binaries, with periods days, following triple evolution for different common envelope efficiencies (). The top, right panels show the orbital period as a function of eccentricity for these systems. The binaries in Figures 3 and 4 are no longer in triples, but the stellar companion in these systems was the tertiary star of the original triple of the time. The initial period and eccentricity of the companion are shown in Figure 5. The bottom panels of Figures 3 and 4 show the distributions for binaries evolved in isolation.
@@
Gaia BH1-like systems form from the initial primary and secondary star, after a common-envelope phase. (We define a Gaia BH1-like system to be one where the eccentricity is within a factor of 1.5 of the observed one, while the BH mass, secondary mass, and period are within a factor of two of those observed). In binaries, the separation before the common envelope only evolves adiabatically due to wind mass loss (with constant eccentricity). Therefore, the common envelope occurs at the beginning of the giant phase. Afterward, the separation increases much faster than the radius due to mass loss from the giant star. In triples, the eccentricity of the secondary can evolve via the von Zeipel–Kozai–Lidov effect (Kozai, 1962; Lidov, 1962). Thus, in some cases, the eccentricity can decrease at the beginning of the giant phase, and then increase again so that the common envelope occurs later when the separation between the low-mass star and black hole progenitor is wider, as illustrated in Figures 6 and 7. Moreover, stellar evolution and mass loss can trigger and change such secular evolution (Perets & Kratter, 2012; Michaely & Perets, 2014). Alternatively, the eccentricity of the secondary can be excited by a triple evolution dynamical instability (TEDI) (Perets & Kratter, 2012; Michaely & Perets, 2014; Toonen et al., 2020; Hamers et al., 2022). In this case, the stellar eccentricity is excited, as the system becomes less hierarchical due to stellar evolution. This type of evolution is shown in Figure 8.333The examples in Figures 7– 9 are from a separate batch of simulations run for yr with high output cadence to ensure well-sampled orbital data.
@@ 4 Rates
We now estimate the formation efficiency of Gaia BH1-like systems for triples and binaries. We ran additional simulations to reduce the uncertainty in our rate estimates. For efficiency, we require the secondary star to be between 0.5 and 2 in the new triple simulations, considering the results in § 3, where all Gaia BH1-like binaries form from the inner binary of the progenitor triple.
@@
The formation efficiencies are summarized in Table 1, for triples and binaries. The efficiency is at least a factor of a few higher for triples than binaries.555For much of the enhancement is from the term in equation (1). The triple efficiency would only be 1.5 times greater otherwise. Overall, up to Gaia BH1-like binaries per solar mass of stars formed.
Previously, Di Carlo et al. (2023) computed formation rates for BH-main sequence (BH-MS) binaries from clusters, and from isolated binary evolution (using COSMIC). They find BH-MS binaries (with periods up to 10 yr666Wider systems would not be detectable by Gaia.) per unit star formation from isolated evolution. To compare with this result, we ran additional binary and triple simulations with no restrictions on the companion masses. The formation efficiency BH-MS binaries (with periods less than 10 yr) in these simulations, is summarized in Table 2. For consistency with Di Carlo et al. (2023) we focus on the case.
The formation efficiency of BH-MS binaries for triples is enhanced by a factor of a few compared to binaries.777The formation efficiency for BH-MS from isolated is an order of magnitude greater than in Di Carlo et al., 2023, mostly due to differences in the initial mass ratio distribution. This suggests field triples are competitive with young star clusters for forming BH-MS binaries. (Di Carlo et al., 2023 find that young star clusters are 4 times more efficient than isolated binaries in producing BH-MS binaries.)
Multiple-type | aaGaia BH1-like binaries formed per massive triple or binary. | bbGaia BH1-like binaries per of stars formed, divided by the triple (or binary fraction) of massive stars. | |
@@
Note. — The Possionian uncertainties are
@@ 4.1 Gaia BH2
El-Badry et al. (2023b) reported the discovery of Gaia BH2: a BH-red giant system with a period of 1277 days and an eccentricity of 0.52. The mass of the black hole is , while the companion is a red giant.
For a common envelope efficiency, (), our triple population synthesis produces 0 (3) Gaia BH2-like systems. These systems form from a merger in the inner binary followed by a common envelope. The primary and secondary then evolve into a black hole and red giant. This suggests Gaia BH2-like systems per star formed. Also, all of our Gaia BH2-like systems are more eccentric than the observed system and have eccentricities 0.7.
@@ 5 Alternatives
Secondly, the interaction between the inner binary stars in a triple may have prevented either star from becoming a giant (Justham et al., 2014). This could lead to a low mass tertiary in orbit around a binary black hole system (El-Badry et al., 2023a). This behavior is also not captured by our models. In our triple synthesis, the only systems with binary black holes and a low mass tertiary have periods yr. Thus, they are much wider than the observed systems.
@@ 6 Discussion and Summary
Recently, a couple of candidate BH-stellar binaries have been identified in Gaia data: Gaia BH1 and Gaia BH2. The former is a Sun-like main sequence star in a 185.6-day orbit around a 9.62 black hole, while the latter is a red giant in a 1277-day orbit around an 8.94 black hole. These systems are difficult to explain with isolated binaries, potentially requiring unusual common envelope evolution (El-Badry et al., 2023a) and/or other scenarios, including formation in clusters.
Here, we consider a triple scenario for the formation of such systems. Since, as known from observations, most black hole progenitors are in triples or higher order multiples (Moe & Di Stefano, 2017), this is, to begin with, a more likely scenario than a binary progenitor.
@@
As a result, systems with periods comparable to Gaia BH1 can be produced for smaller common envelope efficiencies (e.g. for ).
@@
The formation efficiency for BH-main sequence binaries with periods yr is a factor of 3-4 higher for triples than for binaries. Thus, triples provide a more consistent scenario for such systems, without the need for very low-efficiency common-envelope scenarios. Dynamical encounters in a cluster are also not required.
@@
BH-main sequence binaries from triples have a more top-heavy BH mass function than those from binaries, due to mergers within the inner binary. For example, the BH never exceeds for a binary initial condition but exceeds this threshold for of triple initial conditions. Almost all BH - main-sequence binaries with BH masses above have periods days. In isolation, such wide binaries would not give rise to any strong interactions, and would not form X-ray binaries, however, flyby encounters in the field can drive such binaries into interaction (Michaely & Perets, 2016). In clusters such wide binaries would be disrupted (soft binaries), but the BHs could be exchanged into other binaries, and even be part of BH-BH binaries that may merge through GW inspirals. Therefore, the elevated BH masses in these systems can affect the overall mass function of GW sources.
@@ Acknowledgments
We thank Ugo Niccolò Di Carlo and Katelyn Breivik for providing information about their binary population synthesis calculations for the Gaia BH binaries. AG is supported at the Technion by a Zuckerman Fellowship.

### v2

@@ Abstract
Recently, several non-interacting black hole-stellar binaries have been identified in Gaia data. For example, Gaia BH1, where a Sun-like star is in a moderate eccentricity (e=0.44), 185-day orbit around a black hole. This orbit is difficult to explain through binary evolution. The present-day separation suggests the progenitor system would have undergone an episode of common envelope evolution, but a common envelope should shrink the period below the observed one. Since the majority of massive stars form in higher multiplicity systems, a triple evolution scenario is more likely for the progenitors of BH binaries. Here we show that such systems can indeed be more easily explained via evolution in hierarchical triple systems. von Zeipel–Lidov–Kozai oscillations or instabilities can delay the onset of the common envelope phase in the inner binary of the triple, so that the black hole progenitor and low-mass star are more widely separated when it begins, leading to the formation of wider binaries. There are also systems with similar periods but larger eccentricities, where the BH progenitor is a merger product of the inner binary in the triple. Such mergers lead to a more top-heavy black hole mass function.
@@ 1 Introduction
Approximately one out of every thousand stars will end their lives as a black hole (BH). This means a galaxy like the Milky Way should contain of order BHs. However, only of order a hundred BH systems are known in our Galaxy, predominantly from studies of X-ray binaries (Corral-Santana et al., 2016; Fortin et al., 2023), that can only identify close, interacting systems.
Microlensing (Lam et al., 2022; Mróz et al., 2022; Sahu et al., 2022), spectroscopy (Shenar et al., 2022), and astrometry (Andrews et al., 2022; Shahaf et al., 2023) can be used to identify isolated BHs or non-interacting binaries. Such systems probe high mass binary evolution over a wider range of parameter space than X-ray binaries, providing important constraints on the BH natal kick distribution and common envelope physics.
Gaia DR3 contains a few binary orbital solutions (Gaia Collaboration et al., 2023a, b), prompting searches for astrometric BH binaries (Andrews et al., 2022; Shahaf et al., 2023; Chakrabarti et al., 2023; El-Badry et al., 2023a, b). To date, two unambiguous BH-stellar binaries have been identified: Gaia BH1 and Gaia BH2. Gaia BH1 contains a G-type main sequence star in a 185.5-day orbit around a BH. The star’s orbit is moderately eccentric with (Chakrabarti et al., 2023; El-Badry et al., 2023a). Gaia BH2 contains a red giant in a 1276.7 day orbit around a BH. This star’s orbit is also moderately eccentric with (El-Badry et al., 2023b). The uncertainties on the orbital periods and eccentricities are much less than 1%.
These systems pose a challenge to isolated binary evolution models. On the one hand, the present-day separations suggest that the progenitor system would have undergone an episode of common envelope evolution. On the other hand, a common envelope would shrink the binary period below the observed one. The tension can be resolved by invoking an unusually large common envelope efficiency ( for Gaia BH1; El-Badry et al., 2023a), or a very massive progenitor for the BH that avoids a red giant phase (and hence the common envelope) entirely (El-Badry et al., 2023b).
In any case, an isolated binary is not the most likely initial configuration for the progenitor system. At least 55% of O star primaries have two or more companions with mass ratio, (Moe & Di Stefano, 2017). Stellar evolution in triples could give rise to a variety of novel evolutionary channels and produce different types of binaries and merger products (e.g. Perets & Fabrycky, 2009; Perets & Kratter, 2012; Antognini et al., 2014; Michaely & Perets, 2014; Naoz, 2016; Antonini et al., 2017; Rose et al., 2019; Toonen et al., 2020; Stegmann et al., 2022; Hamers et al., 2022; Toonen et al., 2022; Kummer et al., 2023; Shariat et al., 2023). Motivated by this observation we check whether triple evolution can reproduce the Gaia BH binaries. We find that triples can reproduce the observed period and eccentricity of Gaia BH1, even for normal common envelope efficiencies ().
@@ 2.1 Initial conditions
The progenitor of the BH in Gaia BH1 would be , while its companion is 0.93 . Thus, the initial mass ratio is . The distribution of binary properties for such low mass ratios is unconstrained by Moe & Di Stefano (2017). Here, we extrapolate the broken power-law mass ratio distribution from their work down to the brown dwarf boundary. This extrapolation affects the distribution of companion periods, as shown in Figure 1.
@@ 2.2 Triple evolution
MSE does not properly handle supernovae during common envelope phases, and we filter out such systems. This only occurs in a narrow region of parameters space with initially primaries that experience a common envelope during an AGB phase. Such primaries form BHs. This filter does not affect our conclusions on the formation of Gaia BH-like systems or on the overall formation BH binaries.
@@ 3 Results
The top panels of Figure 3 show the initial period and eccentricity of stars that end up in BH-main sequence binaries with periods days, following triple evolution. The top, left panels of Figure 4 and Figure 5 show the final BH masses and eccentricities for these systems. The top, right panels show the final period as a function of eccentricity. The binaries in Figures 4 and 5 are no longer in triples, but the stellar companion in these systems was the tertiary star of the original triple of the time. The bottom panels of Figures 3 and Figures 4 and 5 show the same distributions for binaries evolved in isolation.
@@
Gaia BH1-like systems form from the initial primary and secondary star, after a common-envelope phase. (We define a Gaia BH1-like system to be one where the eccentricity is between 2/3 and 3/2 times the observed one, while the BH mass, secondary mass, and period are between half and twice the observed ones). In binaries, the separation before the common envelope only evolves adiabatically due to wind mass loss (with constant eccentricity). Therefore, the common envelope occurs at the beginning of the giant phase. Afterward, the separation increases much faster than the radius due to mass loss from the giant star. In triples, the eccentricity of the secondary can evolve via the von Zeipel–Kozai–Lidov effect (Kozai, 1962; Lidov, 1962; Naoz, 2016). Thus, in some cases, the eccentricity can decrease at the beginning of the giant phase, and then increase again so that the common envelope occurs later when the separation between the low-mass star and BH progenitor is wider, as illustrated in Figures 6, 7, and 8. Moreover, stellar evolution and mass loss can trigger and change such secular evolution (Perets & Kratter, 2012; Michaely & Perets, 2014). Alternatively, the eccentricity of the secondary can be excited by a triple evolution dynamical instability (TEDI) (Perets & Kratter, 2012; Michaely & Perets, 2014; Toonen et al., 2020; Hamers et al., 2022). In this case, the stellar eccentricity is excited, as the system becomes less hierarchical due to stellar evolution. This type of evolution is shown in Figure 9.333The examples in Figures 8– 10 are from a separate batch of simulations run for yr with high output cadence to ensure well-sampled orbital data.
@@
Rastello et al. (2023), propose an alternative definition of Gaia BH1-like binaries with more stringent requirements on the component masses. In particular, they require the BH to be and the companion to be . With this definition, only one Gaia BH1-like system is produced for .
@@ 4 Rates
We now estimate the formation efficiency of Gaia BH1-like systems for triples and binaries. We ran of order additional triple and of order additional binary simulations to reduce the uncertainty in our rate estimates. For efficiency, we require the secondary star to be between 0.5 and 2 in of the new triple simulations, considering the results in § 3, where all Gaia BH1-like binaries form from the inner binary of the progenitor triple.
@@
The formation efficiencies are summarized in Table 1, for triples and binaries. The efficiency is at least a factor of a few higher for triples than binaries. However, for the enhancement is mostly due to the larger fraction of solar-type stars in our triple initial conditions (the term in equation 1). Otherwise, triples are not more efficient than binaries in producing Gaia BH1-like binaries for . Overall, there are up to Gaia BH1-like binaries per solar mass of stars formed.
Previous work has estimated the formation rate of BH-main sequence binaries in open star clusters (Rastello et al., 2023; Di Carlo et al., 2023; Tanikawa et al., 2024). For example, Rastello et al. (2023) find a formation efficiency of for open clusters. This is within a factor of two of the maximum formation efficiency for triples. Other references give different rates (e.g. in Tanikawa et al., 2024), but this is primarily due to difference definitions of Gaia BH1-like binaries.
We now estimate formation efficiency of all BH-MS binaries. Previously, Di Carlo et al. (2023) find a total of () BH-MS binaries (with periods up to 10 yr555Wider systems would not be detectable by Gaia.) per unit star formation from isolated binary evolution (open clusters). Tanikawa et al. (2024) obtained similar results. To compare with these results, we ran additional binary and triple simulations with no restrictions on the companion masses. The formation efficiency BH-MS binaries (with periods less than 10 yr) in these simulations, is summarized in Table 2.
The formation efficiency of BH-MS binaries for triples is enhanced by a factor of a few compared to binaries.666The formation efficiency for BH-MS from isolated evolution is an order of magnitude greater than in Di Carlo et al. 2023, mostly due to differences in the initial mass ratio distribution. This enhancement is mostly due to the larger fraction of low mass companions in the triple initial conditions, compared to binaries. Overall, field triples are competitive with open clusters for forming BH-MS binaries.
| Multiple-type | aaGaia BH1-like binaries formed per massive triple or binary. The range shows the uncertainty from different definitions of Gaia BH1-like binaries. The smaller efficiency corresponds to the definition in Rastello et al. (2023). For the larger efficiency, a Gaia BH1-like system is one where the eccentricity is within a factor of 1.5 of the observed system, while the BH mass, secondary mass, and period are within a factor of two of the observed system. | bbGaia BH1-like binaries per of stars formed, divided by the triple (or binary fraction) of massive stars. | |
@@
Note. — The Possonian uncertainties are , except for the second row. There the rate estimate is based on a single BH1-like system, and thus the uncertainty is of order unity.
@@ 4.1 Gaia BH2
El-Badry et al. (2023b) reported the discovery of Gaia BH2: a BH-red giant system with a period of 1277 days and an eccentricity of 0.52. The mass of the BH is , while the companion is a red giant.
For a common envelope efficiency, (), our triple population synthesis produces 0 (3) Gaia BH2-like systems. These systems form from a merger in the inner binary followed by a common envelope. The primary and secondary then evolve into a BH and red giant. This suggests Gaia BH2-like systems per star formed. Also, all of our Gaia BH2-like systems are more eccentric than the observed system and have eccentricities 0.7. For isolated binaries, there are no Gaia BH2-like binaries for , while for the formation efficiency is an order of magnitude smaller than for triples.
@@ 5 Alternatives
Secondly, the interaction between the inner binary stars in a triple may have prevented either star from becoming a giant (Justham et al., 2014). This could lead to a low mass tertiary in orbit around a binary BH system (El-Badry et al., 2023a). This behavior is also not captured by our models. In our triple synthesis, the only systems with binary BHs and a low mass tertiary have periods days. Thus, they are much wider than the observed systems.
@@ 6 Discussion and Summary
Recently, a couple of candidate BH-stellar binaries have been identified in Gaia data: Gaia BH1 and Gaia BH2. The former is a Sun-like main sequence star in a 185.5-day orbit around a 9.32 BH, while the latter is a red giant in a 1277-day orbit around an 8.94 BH. These systems are difficult to explain with isolated binaries, potentially requiring unusual common envelope evolution (El-Badry et al., 2023a) and/or other scenarios, including formation in clusters.
Here, we consider a triple scenario for the formation of such systems. Since, as known from observations, most BH progenitors are in triples or higher order multiples (Moe & Di Stefano, 2017), this is, to begin with, a more likely scenario than a binary progenitor.
@@
As a result, systems with periods comparable to Gaia BH1 can be produced for smaller common envelope efficiencies (e.g. for ), though there will only be a few per solar mass. For both binaries and triples can produce Gaia BH1-like systems. The triple formation efficiency () is a factor of a few higher, due to the larger fraction of low mass companions in the triple initial conditions. For the formation efficiency of Gaia BH1-like systems in triples is comparable to literature estimates for open clusters. Furthermore, triple dynamics in clusters may enhance the formation efficiency in these environments. We leave this to future study.
@@
The formation efficiency for BH-main sequence binaries with periods yr is a factor of higher for triples than for binaries. Once again, this is most likely due to the larger fraction of low mass companions in our triple initial conditions, rather than differences in dynamics. The maximum formation efficiency of BH-MS systems in the triple channel is competitive with open clusters.
@@
BH-main sequence binaries from triples have a more top-heavy BH mass function than those from binaries, due to mergers within the inner binary. For example, the BH never exceeds for a binary initial condition but exceeds this threshold for of triple initial conditions. Almost all BH - main-sequence binaries with BH masses above have periods days. In isolation, such wide binaries would not give rise to any strong interactions, and would not form X-ray binaries, however, flyby encounters in the field can drive such binaries into interaction (Michaely & Perets, 2016). In clusters such wide binaries would be disrupted (soft binaries), but the BHs could be exchanged into other binaries, and even be part of BH-BH binaries that may merge through GW inspirals. Therefore, the elevated BH masses in these systems can affect the overall mass function of GW sources. In principle, flybys may also perturb triple phases of evolution (Michaely & Perets, 2019, 2020; Shariat et al., 2023). However, in our case most triples are destroyed on timescales comparable to the lifetime of the massive primary.
@@ Acknowledgments
We thank the anonymous referee for a constructive report. We thank Sara Rastello, Ataru Tanikawa, and Sukanya Chakrabarti for suggesting useful references. We thank Ugo Niccolò Di Carlo and Katelyn Breivik for providing information about their binary population synthesis calculations for the Gaia BH binaries. AG is supported at the Technion by a Zuckerman Fellowship.